# DOMBA: Double Model Balancing for Access-Controlled Language Models via Minimum-Bounded Aggregation

### v1

@@ Abstract
The utility of large language models (LLMs) depends heavily on the quality and quantity of their training data. Many organizations possess large data corpora that could be leveraged to train or fine-tune LLMs tailored to their specific needs. However, these datasets often come with access restrictions that are based on user privileges and enforced by access control mechanisms. Training LLMs on such datasets could result in exposure of sensitive information to unauthorized users. A straightforward approach for preventing such exposure is to train a separate model for each access level. This, however, may result in low utility models due to the limited amount of training data per model compared to the amount in the entire organizational corpus. Another approach is to train a single LLM on all the data while limiting the exposure of unauthorized information. However, current exposure-limiting methods for LLMs are ineffective for access-controlled data, where sensitive information appears frequently across many training examples. We propose DOMBA– double model balancing – a simple approach for training and deploying LLMs that provides high utility and access-control functionality with security guarantees. DOMBA aggregates the probability distributions of two models, each trained on documents with (potentially many) different access levels, using a “min-bounded" average function (a function that is bounded by the smaller value, e.g., harmonic mean). A detailed mathematical analysis and extensive evaluation show that DOMBA safeguards restricted information while offering utility comparable to non-secure models.
@@ Introduction
Figure 1: Two main phases of our proposed method. (a) Training: Documents are grouped and divided into two partitions according to their access levels (each access level is depicted in a different color). Training includes two steps: DOMBA-INIT: A submodel is trained on each partition, resulting in $M_{1}$ and $M_{2}$. DOMBA-FT: To achieve a model dedicated to access level B (AL-B), $M_{2}$ is fine-tuned on AL-B documents, resulting in $M_{2}^{B}$ (note that AL-B documents are included in the training sets of both $M_{1}$ and $M_{2}^{B}$). (b) Inference: Given a context and access level, the corresponding submodels are selected, and their probability distributions are aggregated using a “min-bounded" aggregation function $f_{MB}$.
@@
In this paper, we propose DOMBA, a method for training and deploying LLMs that incorporates an access-control mechanism while maintaining high utility. An overview of DOMBA is presented in Figure 1. To protect sensitive information, DOMBA “balances" two submodels (trained on two different data partitions, each including different access levels) during inference, using a min-bounded average function; intuitively, each submodel “knows" different restricted information. During text generation, the min-bounded function makes it unlikely for information known to just one submodel to be generated. Assuming that restricted information is not shared between the two partitions, this ensures that no restricted information is likely to be generated. We note that it is important that the division of access levels into partitions is done such that access levels with shared sensitive information are assigned to the same partition.
@@
To analyze the privacy protection provided by DOMBA, we formalize the notion of “exposure of secrets." This notion is based on the change in probability of a token (relative to other tokens) between two language models. The greater the change in a token’s probability (relative to other tokens) when using one model over the other, the greater the exposure of that token. In our approach, the exposure of DOMBA over both submodels is bounded by the best possible value (i.e., replacing DOMBA by any other model would not improve the bound). This means that DOMBA limits the exposure of “secrets" that appear in just one partition (since one of the submodels does not “know" them).
@@
Evaluating sensitive information exposure is challenging due to the fact that defining what constitutes sensitive information is a complex and context-dependent task, varying based on organizational policy and other factors (Brown et al. 2022). Measuring average-case sensitive information exposure is inadequate, since an adversary could devise a prompt that causes the model to substantially deviate from average behavior (Wolf et al. 2024). To address these challenges, we introduce three new security metrics for evaluating sensitive information exposure. The first metric assesses worst-case and “extreme-case" exposure across data records. In the second metric, we evaluate the probabilities assigned to certain tokens that should not be exposed by a secure model. The third new metric involves a “secret inference attack" (SIA) that is based on membership inference attack (MIA) techniques. In addition to evaluating DOMBA on the three new metrics, we also evaluate it using the canary technique (Carlini et al. 2019), in which specific phrases (canaries) are injected into the training data, and the model’s inclination to generate them compared to similar phrases not included in the training data is measured.
@@
    Developing a mathematical framework to model language models’ exposure of sensitive information and establishing bounds for DOMBA’s exposure.
@@
Very few studies have addressed the use of LLMs in the access-control scenario: Tiwari et al. (2023) proposed using mixture of experts (MoE) in conjunction with training a separate model for each access level in order to support users with multiple access rights. However, this approach relies solely on non-restricted documents, which substantially reduces its utility. Wutschitz et al. (2023) proposed using retrieval-augmented generation (RAG) with access rights, which prevents the retrieval of unauthorized documents. As illustrated by Tiwari et al. (2023), using RAG by itself, without training the model on the access-controlled data, may be insufficient to achieve a substantial adjustment of LLM behaviors, such as altering writing styles or tones and incorporating new domain-specific knowledge and terminology. Therefore, securely using LLMs trained on access-controlled data, which is the purpose of our study, is essential.
Several studies have explored the use of differential private training algorithms for deep learning (Yu et al. 2022; Abadi et al. 2016). These algorithms are designed to provide privacy guarantees when each “secret" appears in only one or a few data records. Scaling these algorithms to protect privacy when hundreds of documents contain the same sensitive information results in a substantial degradation in utility (Ginart et al. 2022). Differencial privacy (DP) federated learning (Geyer, Klein, and Nabi 2018) is a specific application of DP in deep learning in which contributions from different clients are aggregated during training, and noise is added to maintain DP. While federated learning might seem promising for the access-control scenario (with each access level treated as a client), Wei et al. (2020) demonstrated that performance drops substantially when using a small amount of clients (access levels). In contrast, DOMBA inherently does not depend on the number of access levels.
@@
In addition to DP-based approaches, sanitization methods have been proposed to empirically reduce sensitive information exposure by anonymizing names, numbers, etc. (Lison et al. 2021). Sanitization may be insufficient for the protection of organizational data which can include sensitive text that does not contain a specific name or number (e.g., “the product performs worse than expected" may be considered sensitive text).
One might consider using alignment methods for protecting sensitive information. Such methods include prompt engineering (Chen et al. 2023) and reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022), which is based on instructing the model to behave in a certain way or fine-tuning it to alter its behavior. These methods have been shown to be vulnerable to data extraction attacks and “jailbreaking" (Liu et al. 2024; Chen et al. 2023; Nasr et al. 2023); in addition, theoretical work by Wolf et al. (2024) suggests that an adversary can bypass model alignment methods with a long enough query.
Another approach for mitigating sensitive information exposure is using a privacy regularized loss function (Mireshghallah et al. 2021), however this approach does not provide theoretical guarantees. Unlike previous methods, DOMBA addresses the problem by directly managing probability distributions without relying on the sensitive data type or depending on regularization.
@@
Exposure of Secrets
We begin by defining exposing a secret in a formal sense. As highlighted by Brown et al. (2022), secrecy is relative – something is deemed secret if it is known by some but unknown to others. Therefore, our concept of secrecy involves comparing probabilities assigned to a token by two models. One possible approach is to use the ratio of the probabilities assigned by the models to assess secrecy. However, this method has drawbacks. Consider the following probability distributions over the tokens $a,b,c,d$: $p_{1}=(0.7,0.1,0.1,0.1)$ and $p_{2}=(0.97,0.01,0.01,0.01)$. The probabilities’ ratios ($p_{1}/p_{2}$) are $(0.72,10,10,10)$. This implies that tokens $b,c$, and $d$ are “secret" in $p_{1}$ compared to $p_{2}$. However, it seems more appropriate to consider $a$ as secret in $p_{2}$ compared to $p_{1}$, because $p_{2}$ assigns $a$ a probability that is 97 higher than all other tokens, whereas $p_{1}$ assigns it a probability that is only seven times higher. To address this, we compare the probability ratio of a token $t$ (between two models) to a “typical probability ratio" (TPR).
@@
Let $c$ be a context, and let $M_{1},M_{2}$ be language models. We define the “TPR at $c$" of $M_{1},M_{2}$ as $tpr_{c}(M_{1},M_{2})=GM(\frac{p_{M_{1}}(t|c)}{p_{M_{2}}(t|c)})$.
@@
Exposure Properties
In this subsection, we explore certain properties of exposure that are essential for later discussions.
@@
Let $c$ be a context, and let $M$ be a language model. We define the “typical probability at $c$" of $M$ as $tp_{c}(M)=GM(p_{M}(t|c))$.
Let $t$ be a token, we further define the “relative probability of $t$ at $c$ by $M$" as $rp_{M}(t|c):=\frac{p_{M}(t|c)}{tp_{c}(M)}$.
@@
Lemma 2 (Exposure-multiplicity).
@@
    $\alpha=-1$ (Harmonic mean): $f(x,y)=\frac{2}{x}+\frac{2}{y},\lambda_{f}=2$.
@@
While it might be unclear how to interpret $\overline{f_{c}}(M_{1},M_{2})$, it is related to a notion of “mean exposure" between $M_{1}$ and $M_{2}$:
@@
Let $c$ be a context and $M_{1},M_{2}$ be language models. We define the “mean absolute exposure between $M_{1}$ and $M_{2}$ at $c$" as $MAE_{c}(M_{1},M_{2})=GM(\max(\frac{rp_{M_{1}}(t|c)}{rp_{M_{2}}(t|c)},\frac{rp_{M_{2}}(t|c)}{rp_{M_{1}}(t|c)}))$.
@@
Since access-controlled datasets are not publicly available, we required datasets that mimic the access-control scenario. These datasets need to be divided into different topics (which serve as access levels), with many data records per topic. Additionally, to use two of our security evaluation metrics, the data records should contain phrases that we refer to as “sensitive-mimicking phrases" – phrases unique to the topic that could be considered sensitive / secret.
@@
For training we used LORA (Hu et al. 2021). LORA is a fine-tuning technique that uses a small number of trainable parameters. Training is relatively fast with LORA, and the resulting model requires minimal storage space. These qualities were crucial for our experiments, as we conducted numerous trials with limited computational resources. However, it is important to note that the theoretical analysis is not dependent on the training method, and we anticipate that the experiments can be replicated using other training techniques as well. The base model used was OpenAI-GPT (Radford et al. 2018) which has 117 million parameters and a vocabulary size of 40,478. This model’s original training data is a books dataset from 2015 (Zhu et al. 2015). This limits the prior knowledge the model possesses regarding movies and recipes. Since recent LLMs are trained on more recent and diverse datasets, evaluating them on sensitive information from the movie and recipe datasets would be challenging, as the models are probably familiar with some of the information. We note that although recent LLMs are larger and perform better than OpenAI-GPT, many are still based on the same underlying principles. Our theoretical analysis and proposed approach generalizes to any language model based on next-token prediction and does not rely on the specifics of any particular architecture. Regarding training parameters, we conducted experiments with varying numbers of training epochs (1, 2, and 4). The hyperparameters for LORA were set to default values and were not explored: r=64, lora_alpha=32, lora_dropout=0.05, optimizer=paged_adamw_32bit, learning rate=5e-4, and warmup_ratio = 0.03. All experiments were conducted using an NVIDIA A100-SXM4-40GB GPU.
@@
In these models, which serve as baselines, no attempt is made to secure sensitive information. FT-ALL: OpenAI-GPT fine-tuned on the entire training dataset. AGG-A: Similar to DOMBA-INIT, but using arithmetic mean, a non-min-bounded function, during aggregation.
@@
While these models are trained on all the data with an effort made to secure sensitive information, they do not include an access-control mechanism. SUBMIX: An aggregated model constructed using the method of Ginart et al. (2022), with three submodels (two parts + the base model). For a meaningful comparison, we tuned the privacy parameter $\beta$ to 0.3, which resulted in utility comparable to DOMBA on the movies dataset. D-I-H: DOMBA-INIT (without DOMBA-FT), using harmonic mean for aggregation. D-I-M: DOMBA-INIT (without DOMBA-FT), using minimum for aggregation.
@@
Metrics
Utility Evaluation
We evaluate utility in terms of perplexity on two evaluation sets as follows: 1. HOPPL: perplexity on held out data with access levels that were not used for training. This metric provides a “fair" way of comparing secure and non-secure models, as the non-secure models are not expected to gain by “knowing" restricted information. 2. AUPPL: perplexity on held out data of the access levels used for training (for access-controlled models - the corresponding variant is used for each access level). The main purpose of this metric is to compare the utility of secure and access-controlled models. We expect the access-controlled models to gain utility by “knowing" authorized restricted information. For both metrics above, we calculate the perplexity as: $perp_{M}(D_{e})=\exp(\frac{1}{|D_{e}|}\sum_{r\in D_{e}}\sum_{i}{-\log(p_{M}(r_{i}|r_{<i}))}$, where $|D_{e}|$ is the amount of tokens in $D_{e}$, $r$ is a data record, $r_{i}$ is the i’th token in the record, $r_{<i}$ are the tokens preceding it, and $p_{M}$ is the probability assigned by the model.
Exposure (EXP)
In our theoretical analysis (Theorem 2), we established that the exposure of $M=DAGG_{f}(M_{1},M_{2})$ over both $M_{1}$ and $M_{2}$ is bounded for any token by $\lambda_{f}\overline{f_{c}}(M_{1},M_{2})\leq\lambda_{f}\sqrt{MAE_{c}(M_{1},M_{2})}$. To validate this, we measure “extreme case" exposure of $M$ over $M_{1}$ and $M_{2}$. We report the maximum and 99th percentile exposure ($=\frac{rp_{M}(t|c)}{\min(rp_{M_{1}}(t|c),rp_{M_{2}}(t|c))}$) for all tokens observed in the data, given the previous tokens as context.
Secret Perplexity (SPPL)
Secret Inference Attack (SIA)
This attack is based on a membership inference attack with a reference model (Mireshghallah et al. 2022; Kumar Murakonda, Shokri, and Theodorakopoulos 2021). The original attack works as follows: Given a reference model $M_{b}$, a target model $M$, and a potential training data record $r$ of $M$, measure the log ratio of the probabilities of $r$ according to $M$ and $M_{b}$, that is $\log(\frac{p_{M}(r)}{p_{M_{b}}(r)})$. If this value is above a certain threshold, consider $r$ as belonging to the training data of $M$. In our scenario, instead of inferring the membership of any data record, the adversary tries to infer secrets. Therefore, we only consider probabilities assigned to sensitive-mimicking phrases: cast members’ names for the movie review dataset and secret ingredients for the recipe dataset. The attack dataset consists of tuples $(c,t,label)$, where $c$ is a context, $t$ is a phrase, and $label$ is $true$ if $t$ is sensitive and $false$ otherwise. To obtain data points labeled $false$, we replace each sensitive-mimicking phrase $t$ by $t^{\prime}$, which is another phrase of the same type (cast member name or ingredient) that is not a sensitive-mimicking phrase. For every data point $(c,t,true)$, we have a data point $(c,t^{\prime},false)$. We report the AUC-ROC of the attack.
The Canary Technique (CAN)
@@
Table 1: Results with two epochs of training. For each metric and model, the results for both the recipe dataset (R) and movie review dataset (M) are presented. The best values for secure and access-controlled models are in bold, and the second best values are underlined. (Note that PER-AL is trivially secure and therefore, no results are presented for its security; the EXP metric is only meaningful for methods which aggregate two submodels and therefore, it is not presented for FT-ALL.)

### v2

@@ Abstract
The utility of large language models (LLMs) depends heavily on the quality and quantity of their training data. Many organizations possess large data corpora that could be leveraged to train or fine-tune LLMs tailored to their specific needs. However, these datasets often come with access restrictions that are based on user privileges and enforced by access control mechanisms. Training LLMs on such datasets could result in exposure of sensitive information to unauthorized users. A straightforward approach for preventing such exposure is to train a separate model for each access level. This, however, may result in low utility models due to the limited amount of training data per model compared to the amount in the entire organizational corpus. Another approach is to train a single LLM on all the data while limiting the exposure of unauthorized information. However, current exposure-limiting methods for LLMs are ineffective for access-controlled data, where sensitive information appears frequently across many training examples. We propose DOMBA– double model balancing – a simple approach for training and deploying LLMs that provides high utility and access-control functionality with security guarantees. DOMBA aggregates the probability distributions of two models, each trained on documents with (potentially many) different access levels, using a “min-bounded” average function (a function that is bounded by the smaller value, e.g., harmonic mean). A detailed mathematical analysis and extensive evaluation show that DOMBA safeguards restricted information while offering utility comparable to non-secure models.
Code and datasetshttps://github.com/ppo1/DOMBA
@@ Introduction
Figure 1: Two main phases of our proposed method. (a) Training: Documents are grouped and divided into two partitions according to their access levels (each access level is depicted in a different color). Training includes two steps: DOMBA-INIT: A submodel is trained on each partition, resulting in $M_{1}$ and $M_{2}$. DOMBA-FT: To achieve a model dedicated to access level B (AL-B), $M_{2}$ is fine-tuned on AL-B documents, resulting in $M_{2}^{B}$ (note that AL-B documents are included in the training sets of both $M_{1}$ and $M_{2}^{B}$). (b) Inference: Given a context and access level, the corresponding submodels are selected, and their probability distributions are aggregated using a “min-bounded” aggregation function $f_{MB}$.
@@
In this paper, we propose DOMBA, a method for training and deploying LLMs that incorporates an access-control mechanism while maintaining high utility. An overview of DOMBA is presented in Figure 1. To protect sensitive information, DOMBA “balances” two submodels (trained on two different data partitions, each including different access levels) during inference, using a min-bounded average function; intuitively, each submodel “knows” different restricted information. During text generation, the min-bounded function makes it unlikely for information known to just one submodel to be generated. Assuming that restricted information is not shared between the two partitions, this ensures that no restricted information is likely to be generated. It is crucial to separate access levels into partitions in a way that groups access levels with shared sensitive information into the same partition. If we represent access levels as nodes in a graph and connect nodes with edges whenever they share sensitive information, such separation is achievable if the graph is disconnected (and, for practical purposes, does not contain an overly large connected component).
@@
To analyze the privacy protection provided by DOMBA, we formalize the notion of “exposure of secrets” by developing a framework called ”token exposure.” Token exposure is based on the change in probability of a token (relative to other tokens) between two language models. The greater the change in a token’s probability (relative to other tokens) when using one model over the other, the greater the exposure of that token. Using our approach, the exposure of DOMBA over both submodels is bounded by the best possible value (i.e., replacing DOMBA by any other model would not improve the bound). This means that DOMBA limits the exposure of “secrets” that appear in just one partition (since one of the submodels does not “know” them).
@@
Evaluating sensitive information exposure is challenging due to the fact that defining what constitutes sensitive information is a complex and context-dependent task, varying based on organizational policy and other factors (Brown et al. 2022). Measuring average-case sensitive information exposure is inadequate, since an adversary could devise a prompt that causes the model to substantially deviate from average behavior (Wolf et al. 2024). To address these challenges, we introduce three new security metrics for evaluating sensitive information exposure. The first metric assesses worst-case and “extreme-case” exposure across data records. In the second metric, we evaluate the probabilities assigned to certain tokens that should not be exposed by a secure model. The third new metric involves a “secret inference attack” (SIA) that is based on membership inference attack (MIA) techniques. In addition to evaluating DOMBA on the three new metrics, we also evaluate it using the canary technique (Carlini et al. 2019), in which specific phrases (canaries) are injected into the training data, and the model’s inclination to generate them compared to similar phrases not included in the training data is measured.
@@
    Developing the mathematical framework of token exposure to analyse language models’ exposure of sensitive information and establishing bounds for DOMBA.
@@
Very few studies have addressed the use of LLMs in the access-control scenario: Tiwari et al. (2023) proposed using mixture of experts (MoE) in conjunction with training a separate model for each access level in order to support users with multiple access rights. However, this approach relies solely on non-restricted documents, which substantially reduces its utility. Wutschitz et al. (2023) proposed using retrieval-augmented generation (RAG) with access rights, which prevents the retrieval of unauthorized documents. As illustrated by Tiwari et al. (2023), using RAG by itself, without training the model on the data, may be insufficient to achieve a substantial adjustment of LLM behaviors, such as altering writing styles or tones and incorporating new domain-specific knowledge and terminology. We consider RAG to be complementary to DOMBA, with the choice between them, or a combination of both, being highly dependent on the specific task and dataset. Determining the optimal approach for each case is beyond the scope of this work.
Several studies have explored the use of differential private training algorithms for deep learning (Yu et al. 2022; Abadi et al. 2016). These algorithms are designed to provide privacy guarantees when each “secret” appears in only one or a few data records. As demonstrated by Ginart et al. (2022), scaling these algorithms to safeguard a partition of sensitive documents on the same topic (which is critical in the access control scenario) results in impractical utility. Differential privacy (DP) federated learning (Geyer, Klein, and Nabi 2018) is a specific application of DP in deep learning in which contributions from different clients are aggregated during training, and noise is added to maintain DP. While federated learning might seem promising for the access-control scenario (with each access level treated as a client), Wei et al. (2020) demonstrated that performance drops substantially when using a small amount of clients (access levels). In contrast, DOMBA inherently does not depend on the number of access levels.
@@
Beyond DP-based approaches, various techniques have been proposed to empirically mitigate the exposure of sensitive information. Such approaches include sanitization (Lison et al. 2021), prompt engineering (Chen et al. 2023), reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022) and using a privacy regularized loss function (Mireshghallah et al. 2021). However, these techniques lack theoretical guarantees, and some have been demonstrated to be vulnerable to attacks such as data extraction and “jailbreaking”  (Liu et al. 2024; Chen et al. 2023; Nasr et al. 2023; Wolf et al. 2024).
@@
Token Exposure
We begin by defining exposing a secret in a formal sense. As highlighted by Brown et al. (2022), secrecy is relative – something is deemed secret if it is known by some but unknown to others. Therefore, our concept of secrecy involves comparing probabilities assigned to a token by two models. One possible approach is to use the ratio of the probabilities assigned by the models to assess secrecy. However, this method has drawbacks. Consider the following probability distributions over the tokens $a,b,c,d$: $p_{1}=(0.7,0.1,0.1,0.1)$ and $p_{2}=(0.97,0.01,0.01,0.01)$. The probabilities’ ratios ($p_{1}/p_{2}$) are $(0.72,10,10,10)$. This implies that tokens $b,c$, and $d$ are “secret” in $p_{1}$ compared to $p_{2}$. However, it seems more appropriate to consider $a$ as secret in $p_{2}$ compared to $p_{1}$, because $p_{2}$ assigns $a$ a probability that is 97 higher than all other tokens, whereas $p_{1}$ assigns it a probability that is only seven times higher. To address this, we compare the probability ratio of a token $t$ (between two models) to a “typical probability ratio” (TPR).
@@
Let $c$ be a context, and let $M_{1},M_{2}$ be language models. We define the “TPR at $c$” of $M_{1},M_{2}$ as $tpr_{c}(M_{1},M_{2})=GM(\frac{p_{M_{1}}(t|c)}{p_{M_{2}}(t|c)})$.
@@
Token Exposure Properties
In this subsection, we explore certain properties of token exposure that are essential for later discussions.
@@
Let $c$ be a context, and let $M$ be a language model. We define the “typical probability at $c$” of $M$ as $tp_{c}(M)=GM(p_{M}(t|c))$.
Let $t$ be a token, we further define the “relative probability of $t$ at $c$ by $M$” as $rp_{M}(t|c):=\frac{p_{M}(t|c)}{tp_{c}(M)}$.
@@
Lemma 2 (Token exposure multiplicity).
@@
    $\alpha=-1$ (Harmonic mean): $f(x,y)=\frac{2xy}{x+y},\lambda_{f}=2$.
@@
While it might be unclear how to interpret $\overline{f_{c}}(M_{1},M_{2})$, it relates to a notion of “mean exposure” between $M_{1}$ and $M_{2}$:
@@
Let $c$ be a context and $M_{1},M_{2}$ be language models. We define the “mean absolute exposure between $M_{1}$ and $M_{2}$ at $c$” as $MAE_{c}(M_{1},M_{2})=GM(\max(\frac{rp_{M_{1}}(t|c)}{rp_{M_{2}}(t|c)},\frac{rp_{M_{2}}(t|c)}{rp_{M_{1}}(t|c)}))$.
@@
Since access-controlled datasets are not publicly available, we required datasets that mimic the access-control scenario. These datasets need to be divided into different topics (which serve as access levels), with many data records per topic. Additionally, to use two of our security evaluation metrics, the data records should contain phrases that we refer to as “sensitive-mimicking phrases” – phrases unique to the topic that could be considered sensitive / secret.
@@
For training we used LORA (Hu et al. 2021). LORA is a fine-tuning technique that uses a small number of trainable parameters. Training is relatively fast with LORA, and the resulting model requires minimal storage space. These qualities were crucial for our experiments, as we conducted numerous trials with limited computational resources. However, it is important to note that the theoretical analysis is not dependent on the training method, and we anticipate that the experiments can be replicated using other training techniques as well. The base model used was OpenAI-GPT (Radford et al. 2018) which has 117 million parameters and a vocabulary size of 40,478. This model’s original training data is a books dataset from 2015 (Zhu et al. 2015). This limits the prior knowledge the model possesses regarding movies and recipes. Since recent LLMs are trained on more recent and diverse datasets, evaluating them on sensitive information from the movie and recipe datasets would be challenging, as the models are probably familiar with some of the information. We note that although recent LLMs are larger and perform better than OpenAI-GPT, many are still based on the same underlying principles. Our theoretical analysis and proposed approach generalizes to any language model based on next-token prediction and does not rely on the specifics of any particular architecture. Regarding training parameters, we conducted experiments with varying numbers of training epochs (1, 2, and 4). The hyperparameters for LORA were set to default values and were not explored: r=64, lora_alpha=32, lora_dropout=0.05, optimizer=paged_adamw_32bit, learning rate=5e-4, and warmup_ratio=0.03. All experiments were conducted on an NVIDIA A100-SXM4-40GB GPU.
@@
In these models, which serve as baselines, no attempt is made to secure sensitive information. FT-ALL: OpenAI-GPT fine-tuned on the entire training dataset. AGG-A: Similar to DOMBA-INIT, but using arithmetic mean, a non-min-bounded function.
@@
While these models are trained on all the data with an effort made to secure sensitive information, they do not include an access-control mechanism. SUBMIX: A DP aggregated model constructed using the method of Ginart et al. (2022), with three submodels (two parts + the base model). For a meaningful comparison, we tuned the privacy parameter $\beta$ to 0.3, which resulted in utility comparable to DOMBA on the movies dataset. D-I-H: DOMBA-INIT (without DOMBA-FT), using harmonic mean for aggregation. D-I-M: DOMBA-INIT (without DOMBA-FT), using minimum for aggregation.
@@
Table 1: Results with two epochs of training. For each metric and model, the results for both the recipe dataset (R) and movie reviews dataset (M) are presented. The best values for secure and access-controlled models are in bold, and the second best values are underlined. (Note that PER-AL is trivially secure and therefore, no results are presented for its security; the EXP metric is only meaningful for methods which aggregate two submodels and therefore, it is not presented for FT-ALL.)
Metrics
Utility Evaluation
We evaluate utility in terms of perplexity on two evaluation sets as follows: 1. HOPPL: perplexity on held out data with access levels that were not used for training. This metric provides a “fair” way of comparing secure and non-secure models, as the non-secure models are not expected to gain by “knowing” restricted information. 2. AUPPL: perplexity on held out data of the access levels used for training (for access-controlled models - the corresponding variant is used for each access level). The main purpose of this metric is to compare the utility of secure and access-controlled models. We expect the access-controlled models to gain utility by “knowing” authorized restricted information. For both metrics above, we calculate the perplexity as: $perp_{M}(D_{e})=\exp(\frac{1}{|D_{e}|}\sum_{r\in D_{e}}\sum_{i}{-\log(p_{M}(r_{i}|r_{<i}))}$, where $|D_{e}|$ is the amount of tokens in $D_{e}$, $r$ is a data record, $r_{i}$ is the i’th token in the record, $r_{<i}$ are the tokens preceding it, and $p_{M}$ is the probability assigned by the model.
Exposure (EXP)
In our theoretical analysis (Theorem 2), we established that the exposure of $M=DAGG_{f}(M_{1},M_{2})$ over both $M_{1}$ and $M_{2}$ is bounded for any token by $\lambda_{f}\overline{f_{c}}(M_{1},M_{2})\leq\lambda_{f}\sqrt{MAE_{c}(M_{1},M_{2})}$. To validate this, we measure “extreme case” exposure of $M$ over $M_{1}$ and $M_{2}$. We report the maximum and 99th percentile exposure ($=\frac{rp_{M}(t|c)}{\min(rp_{M_{1}}(t|c),rp_{M_{2}}(t|c))}$) for all tokens observed in the data, given the previous tokens as context.
Secret Perplexity (SPPL)
Secret Inference Attack (SIA)
This attack is based on a membership inference attack with a reference model (Mireshghallah et al. 2022; Kumar Murakonda, Shokri, and Theodorakopoulos 2021). The original attack works as follows: Given a reference model $M_{b}$, a target model $M$, and a potential training data record $r$ of $M$, measure the log ratio of the probabilities of $r$ according to $M$ and $M_{b}$, that is $\log(\frac{p_{M}(r)}{p_{M_{b}}(r)})$. If this value is above a certain threshold, consider $r$ as belonging to the training data of $M$. In our scenario, instead of inferring the membership of any data record, the adversary tries to infer secrets. Therefore, we only consider probabilities assigned to sensitive-mimicking phrases: cast members’ names for the movie reviews dataset and secret ingredients for the recipe dataset. The attack dataset consists of tuples $(c,t,label)$, where $c$ is a context, $t$ is a phrase, and $label$ is $true$ if $t$ is sensitive and $false$ otherwise. To obtain data points labeled $false$, we replace each sensitive-mimicking phrase $t$ by $t^{\prime}$, which is another phrase of the same type (cast member name or ingredient) that is not a sensitive-mimicking phrase. For every data point $(c,t,true)$, we have a data point $(c,t^{\prime},false)$. We report the AUC-ROC of the attack.
The Canary Technique (CAN)